# TabiBERT:A Large-Scale ModernBERT Foundation Model and Unified Benchmarking Framework for Turkish

### v1

@@ 2.3 Multilingual BERT Models
The development of multilingual pre-trained language models began with Multilingual BERT (mBERT) (Devlin2019BERT), which extended the original BERT architecture to 104 languages by pre-training on a corpus obtained by concatenation of Wikipedia corpora. Despite lacking explicit cross-lingual objectives, mBERT demonstrated surprising zero-shot cross-lingual transfer capabilities (Pires2019MultilingualBERT). XLM (Conneau2019Cross) worked on enhancing cross-lingual alignment and introduced Translation Language Modeling (TLM) that leverages parallel corpora. XLM-RoBERTa (Conneau2020XLM-R) further scaled multilingual pre-training to 2.52.5TB of CommonCrawl data across 100 languages, achieving state-of-the-art performance using only masked language modeling without parallel data. Later studies proposed different pre-training approaches. InfoXLM (Chi2021InfoXLM) applied cross-lingual contrastive learning, ERNIE-M (Ouyang2021ERNIE) used back-translation masked language modeling to create pseudo-parallel data, and Unicoder (Huang2019Unicoder) combined several cross-lingual pre-training objectives. Architectural improvements have also been explored in multilingual encoder models. RemBERT (Chung2020Rethinking) decoupled input and output embeddings in order to better use the model capacity. mDeBERTa (He2023DeBERTaV3) extended DeBERTa’s disentangled attention to the multilingual setting and combined it with ELECTRA-style pre-training, resulting in consistent improvements over XLM-RoBERTa on cross-lingual benchmarks. More recently, XLM-V (Liang2023XLMV) addressed the vocabulary bottleneck by improving token allocation across languages with varying lexical overlap.
@@ 2.3.1 Curse of Multilinguality
A key challenge in multilingual modeling is the curse of multilinguality, a phenomenon where adding more languages to a fixed-capacity model eventually degrades per-language performance (Conneau2020XLM-R). It is demonstrated (Wang2020Negative) that this stems from negative interference between languages competing for shared model parameters, affecting both high-resource and low-resource languages. In a comprehensive study conducted over 10,000 models and across 250 languages (Chang2024Multilinguality), it is observed that while moderate multilinguality can benefit low-resource languages, particularly when linguistically similar languages are added, performance degrades for all languages as the number of languages increases, suggesting that massively multilingual pre-training may not be optimal for any language involved.
@@ 2.4 ModernBERT
ModernBERT is one of the recent and most significant bidirectional encoder-only Transformer models. It upgrades the original BERT architecture by utilizing contemporary enhancements to support longer context lengths and improve efficiency and performance in natural language understanding tasks. It was pre-trained on a massive corpus of  2 trillion tokens of English text and code, enabling stronger generalization and suitability for both language and code understanding. ModernBERT introduces several architectural techniques aimed at improving efficiency, including Rotary Positional Embeddings (RoPE) Su2024Roformer, Flash Attention (Dao2022FlashAttention), unpadding (Zeng2022Boosting), GeGLU activations (Shazeer2020Glu), and alternating attention (Team2024Gemma). Together, these design choices reduce memory usage and inference time compared to conventional encoder models, while preserving strong performance on downstream tasks. These features make ModernBERT effective for classification, retrieval, semantic search, and other NLU tasks, and it serves as an efficient alternative to older BERT variants in research and production settings.
@@ 2.4.1 Language-Specific Variants
llm-jp-modernbert is a similar ModernBERT-based model trained on a large-scale Japanese corpus with $8,192$ token context length (Sugiura2025ModernBERTJapanese). While achieving competitive results on fill-mask evaluations, the model highlighted the importance of high-quality, language-specific pre-training data for optimal performance in downstream tasks.
2.4.2 Domain-Specific Applications
The ModernBERT architecture has proven particularly effective for specialized domains that require robust long-context understanding and efficient representation learning. Clinical ModernBERT was developed as a transformer-based encoder pre-trained on biomedical literature, clinical notes, and medical ontologies, incorporating large-scale resources such as PubMed abstracts and MIMIC-IV clinical data (Lee2025ClinicalModernBERT). Building on ModernBERT’s architectural upgrades, most notably long-context support and efficient attention mechanisms, this domain-adapted model demonstrated strong performance in producing semantically rich representations for long-document biomedical tasks, including clinical classification, retrieval, and entity recognition.
BioClinical ModernBERT further extended this line of work through continual pre-training on the largest biomedical and clinical corpus to date, comprising more than 53.5 billion tokens (Sounack2025BioClinicalModernBERT). Unlike earlier clinical encoders trained on narrow or single-institution datasets, this model uses data from multiple institutions, domains, and regions. This broader training data improves generalization and enables the model to achieve state-of-the-art performance across many biomedical NLP benchmarks.
Beyond English biomedical text, ModernBERT has also been shown to be effective in non-English and highly technical clinical settings. In Japanese radiology report classification, ModernBERT achieved 39% reduction in training time and up to a 1.651.65× speed-up during training and inference, while maintaining comparable classification performance (Yamagishi2025ModernBERTRadiology).
@@ 2.4.3 Architectural Analysis and Comparisons
In a controlled study designed to isolate architectural effects from differences in pre-training data, ModernBERT was compared with DeBERTaV3 when both models were pre-trained on the same dataset. The results indicate that while ModernBERT offers faster training and inference speeds attributable to its architectural optimizations, the earlier DeBERTaV3 architecture maintained superior sample efficiency and overall benchmark performance under identical pre-training conditions (Antoun2025ModernBERT). This finding underscores the importance of disentangling architectural improvements from data effects when evaluating advances in encoder design and suggests that modern architectural enhancements may need to be paired with training methodology innovations to fully realize performance gains.
Another line of architectural comparison came from NeoBERT (Breton2025NeoBERT), a next-generation bidirectional encoder that incorporates state-of-the-art architectural refinements, modern pre-training regimes, and optimized depth-to-width design. NeoBERT achieves state-of-the-art results on large evaluation suites such as GLUE and the Massive Text Embedding Benchmark (MTEB), regardless of its relatively small model size of 250250M parameters and extended context length. It outperforms standard baselines, including BERT Large, RoBERTa Large, NomicBERT, and ModernBERT under identical fine-tuning conditions (Breton2025NeoBERT). This comparison highlights how architectural choices such as optimal layer configurations, positional encoding mechanisms, and more balanced model scaling can yield efficiency and performance advantages even compared to models with greater nominal capacity.
@@ 2.5 Turkish NLP Benchmarks
In response to these limitations, several Turkish NLP benchmark efforts have emerged that aim to consolidate multiple tasks under unified evaluation frameworks. Mukayese aggregated diverse Turkish NLP tasks, including language modeling, machine translation, NER, summarization, and text classification into a single benchmark suite, providing common baselines (Safaya2022Mukayese). TrGLUE adapted the GLUE paradigm to Turkish natural language understanding tasks, though its scope remained largely limited to sentence-level NLU (Altinok2025TrGLUE). More recently, Cetvel, a large-scale LLM-oriented evaluation benchmark, covered a broader set of discriminative and generative tasks while emphasizing cultural and linguistic relevance for Turkish (kuisai2024cetvel).
@@
3 Data
3.1 Pre-training Data
We compiled a comprehensive dataset primarily consisting of Turkish monolingual text, supplemented with a small portion of English, coding and math data, to pre-train our model. To address significant size disparities and differences in quality across corpora and ensure balanced representation of different text types, we employed strategic oversampling of smaller but high-quality datasets. After applying the oversampling strategy, our final pre-training corpus consisted of 86.5886.58B tokens, of which 1313% were non-Turkish sources. In the rest of the paper, all references to the size of the pre-training corpus will include the effects of oversampling. The details of each corpus are explained in the following subsections, and the training corpora statistics, as well as the sampling rates, are summarized in Table 2.
@@ 3.1 Pre-training Data
We extracted the tur_Latn subset of the FineWeb-2 (Penedo2025FineWeb2) as the main Turkish web corpus, with a knowledge cutoff date of April 2024. We also included a sample of 5 million English web pages from the FineWeb-2 English dataset to provide English language context during pre-training. Moreover, we included a Turkish Wikipedia dataset comprising articles scraped from Wikipedia’s Turkish-language pages in August 2024. This dataset provides encyclopedic knowledge covering a wide range of topics written in a formal and informative style.
@@
Table 2: Training Corpora Statistics: The complete pre-training corpus consists of 86.5886.58B tokens.
3.2 Fine-tuning Data
For fine-tuning, we introduce TabiBench, a unified evaluation benchmark meticulously crafted for Turkish NLP. TabiBench encompasses 28 datasets spanning 8 major task categories: text classification, token classification (covering named entity recognition and part-of-speech tagging), semantic textual similarity, natural language inference, question answering, information retrieval, code retrieval, and academic domain challenges. Our benchmark brings together carefully curated 21 Turkish datasets alongside 7 newly created machine-translated datasets, specifically for code retrieval and academic domain evaluation. Every dataset included in TabiBench undergoes a systematic quality assessment and is released with standardized train-validation-test splits to ensure reproducibility and enable fair comparisons across models. Further details on our evaluation datasets and the TabiBench framework are provided in Section 4.4.
4 Methodology
This section elaborates the methodology used in this project, including the tokenizer, model, pre-training objectives, fine-tuning and evaluation.
@@ 4.1 Tokenizer
The tokenizer was trained on a dataset mixture that consists of 95% Turkish and 5% English, code, and math—a similar mixture to our model’s pre-training data. This allows the models to perform well on both Turkish and English languages while supporting code and math.
@@
Fertility (Rust2021HowGood) is a common metric to measure the representation efficiency of a tokenizer, indicating the average number of tokens used to represent text. In a controlled setting where training data is assumed to be ample and diverse, vocabulary size is the main parameter that controls fertility. A large vocabulary allows lower fertility values at the cost of increased model size and, hence, computational and memory requirements (Ali2024Tokenizer). Although a lower fertility is desired in language modeling, there is a clear trade-off between vocabulary size and fertility.
@@
In order to measure the representation efficiency of tokenizers used in this work, we measure fertility values by tokenizing 11M randomly selected web pages and report the relative fertility values compared to TabiBERT in Figure 1. It is observed that:
  • YTU-Cosmos-BERT provides the lowest fertility. This is the result of using an uncased tokenizer.
@@
  • TabiBERT and BERTurk tokenizers, both being cased tokenizers, have comparable fertility values, despite TabiBERT tokenizer having a significantly larger vocabulary size. This is due to TabiBERT tokenizer’s pre-tokenization regex, which results in spending more tokens on numbers, punctuations and line-break characters. This is a conscious design choice and a natural trade-off to better represent these symbols.
@@
  • TurkishBERTweet has a similar fertility value to TabiBERT and BERTurk, despite having a vocabulary size that is thrice and twice that of TabiBERT and BERTurk  respectively, and being an uncased tokenizer. We hypothesize that this is because it was trained and optimized for social media texts compared to other tokenizers that were trained on more general corpora.
  • Lastly, despite its large vocabulary of $256,000$ tokens, the mmBERT tokenizer’s fertility is significantly higher than that of Turkish tokenizers. This is due to Turkish being a relatively low-resource language and having a relatively small amount of data compared to other languages in corpora such as FineWeb-2 (Penedo2025FineWeb2).
[…]
Figure 1: Tokenizer fertility values of established Turkish BERT models and the modern multilingual mmBERT, relative to the proposed TabiBERT (represented by the dashed baseline). X axis labels indicate models along with the vocabulary sizes in parentheses. Y axis indicates relative number of tokens (fertility) required to represent Turkish texts compared to TabiBERT. Despite its large vocabulary of $256,000$ tokens, mmBERT suffers a 41.0% increase in fertility compared to TabiBERT in Turkish texts This demonstrates that despite architectural improvements, the lack of language-specific tokenizer optimization in mmBERT results in substantial subword fragmentation compared to TabiBERT.
4.2 Model
@@
Its inference speed is comparable to other models for a standard context length of 512; however, it is $2.65$ times faster than others when processing long contexts of $8,192$ tokens  (Warner2024ModernBERT). Moreover, for both short and long context lengths, it is can process inputs with twice the batch size thanks to its lower GPU memory footprint.
4.3 Pre-training Objectives
For pre-training, we adopt the Masked Language Modeling (MLM) objective as in ModernBERT, with a 30% token masking rate. We utilize the official ModernBERT pre-training codebase and recipe to ensure methodological consistency and reproducibility. The pre-training process is organized into 3 progressive phases, each exposing the model to 850B, 125B, and 25B tokens, respectively. Hence, we scale the ModernBERT pre-training recipe by 50% with the aim of pre-training for 11T tokens in total.
To best utilize the dataset and maximize training throughput, we pre-tokenize the entire pre-training corpus and chunk it into pieces that fit within the context length. We perform optimization using the StableAdamW algorithm(Wortsman2023StableLowPrecision), paired with a trapezoidal learning rate schedule (Xing2018Walk) and 1-sqrt decay (Hagele2024Scaling) for improved stability during large-scale training. The pre-training data is a mixture of samples from the datasets described in Section 3 and consists of 86.5886.58B tokens, of which 13% are non-Turkish. We pre-train our model for $117$ hours on a cluster of 8 x NVIDIA H100 GPUs, resulting in 11T tokens being exposed to the model, that is, $11.83$ epochs over the entire dataset.
4.4 TabiBench Evaluation Framework
4.4.1 Motivation and Design Principles
The evaluation of Turkish encoder models is constrained by several systemic shortcomings in existing benchmarks. It has long relied on small, task-specific datasets that are repeatedly reused across studies, increasing the risk of test set contamination and overfitting to benchmark idiosyncrasies (Kilinc2017; Yang2023Contamination). In addition, many commonly used resources lack explicit probing of Turkish-specific linguistic phenomena such as morphology and agglutination (Oflazer2018), and do not provide fixed, community-accepted train/validation/test splits, undermining reproducibility and comparability. Recent Turkish BERT studies further illustrate this fragmentation, frequently introducing ad-hoc evaluation datasets or custom splits tailored to individual works (Kesgin2023YTU; Najafi2023TurkishBERTweet). Although emerging benchmark efforts improve task consolidation, they remain limited either in task diversity or in evaluation rigor, and meta-analyses reveal persistent issues in dataset quality, annotation consistency, and cultural adequacy (Umutlu2025Evaluating).
These limitations motivate the development of TabiBench, a standardized evaluation framework explicitly tailored to Turkish transformer-encoder models. TabiBench is designed around three core principles: (i) high-quality, carefully validated datasets with clear documentation and linguistically informed construction, (ii) fixed and publicly released evaluation splits to support reproducible and fair model comparison, and (iii) broad task coverage that extends beyond classical classification and sequence labeling to include modern evaluation dimensions such as academic text understanding, retrieval-oriented reasoning, and code-related tasks. By addressing the fragmentation and quality concerns evident in prior Turkish NLP benchmarks, TabiBench aims to enable more reliable, comprehensive, and future-facing evaluation of Turkish encoder models.
  1. 1.

### v3

@@ 2.3 Multilingual BERT Models
The development of multilingual pretrained language models began with Multilingual BERT (mBERT) (Devlin2019BERT), which extended the original BERT architecture to 104 languages by pretraining on a corpus obtained by concatenation of Wikipedia corpora. Despite lacking explicit cross-lingual objectives, mBERT demonstrated surprising zero-shot cross-lingual transfer capabilities (Pires2019MultilingualBERT). XLM (Conneau2019Cross) worked on enhancing cross-lingual alignment and introduced Translation Language Modeling (TLM) that leverages parallel corpora. XLM-RoBERTa (Conneau2020XLM-R) further scaled multilingual pretraining to 2.52.5TB of CommonCrawl data across 100 languages, achieving state-of-the-art performance using only masked language modeling without parallel data. Later studies proposed different pretraining approaches. InfoXLM (Chi2021InfoXLM) applied cross-lingual contrastive learning, ERNIE-M (Ouyang2021ERNIE) used back-translation masked language modeling to create pseudo-parallel data, and Unicoder (Huang2019Unicoder) combined several cross-lingual pretraining objectives. Architectural improvements have also been explored in multilingual encoder models. RemBERT (Chung2020Rethinking) decoupled input and output embeddings in order to better use the model capacity. mDeBERTa (He2023DeBERTaV3) extended DeBERTa’s disentangled attention to the multilingual setting and combined it with ELECTRA-style pretraining, resulting in consistent improvements over XLM-RoBERTa on cross-lingual benchmarks. More recently, XLM-V (Liang2023XLMV) addressed the vocabulary bottleneck by improving token allocation across languages with varying lexical overlap.
@@ 2.3.1 Curse of Multilinguality
A key challenge in multilingual modeling is the curse of multilinguality, a phenomenon where adding more languages to a fixed-capacity model eventually degrades per-language performance (Conneau2020XLM-R). It is demonstrated (Wang2020Negative) that this stems from negative interference between languages competing for shared model parameters, affecting both high-resource and low-resource languages. In a comprehensive study conducted over 10,000 models and across 250 languages (Chang2024Multilinguality), it is observed that while moderate multilinguality can benefit low-resource languages, particularly when linguistically similar languages are added, performance degrades for all languages as the number of languages increases, suggesting that massively multilingual pretraining may not be optimal for any language involved.
@@ 2.4 ModernBERT
ModernBERT is one of the recent and most significant bidirectional encoder-only Transformer models. It upgrades the original BERT architecture by utilizing contemporary enhancements to support longer context lengths and improve efficiency and performance in natural language understanding tasks. It was pretrained on a massive corpus of  2 trillion tokens of English text and code, enabling stronger generalization and suitability for both language and code understanding. ModernBERT introduces several architectural techniques aimed at improving efficiency, including Rotary Positional Embeddings (RoPE) Su2024Roformer, Flash Attention (Dao2022FlashAttention), unpadding (Zeng2022Boosting), GeGLU activations (Shazeer2020Glu), and alternating attention (Team2024Gemma). Together, these design choices reduce memory usage and inference time compared to conventional encoder models, while preserving strong performance on downstream tasks. These features make ModernBERT effective for classification, retrieval, semantic search, and other NLU tasks, and it serves as an efficient alternative to older BERT variants in research and production settings.
@@ 2.4.1 Language-Specific Variants
llm-jp-modernbert is a similar ModernBERT-based model trained on a large-scale Japanese corpus with $8,192$ token context length (Sugiura2025ModernBERTJapanese). While achieving competitive results on fill-mask evaluations, the model highlighted the importance of high-quality, language-specific pretraining data for optimal performance in downstream tasks.
2.4.2 Domain-Specific Variants
The ModernBERT architecture has proven particularly effective for specialized domains that require robust long-context understanding and efficient representation learning. Clinical ModernBERT was developed as a transformer-based encoder pretrained on biomedical literature, clinical notes, and medical ontologies, incorporating large-scale resources such as PubMed abstracts and MIMIC-IV clinical data (Lee2025ClinicalModernBERT). Building on ModernBERT’s architectural upgrades, most notably long-context support and efficient attention mechanisms, this domain-adapted model demonstrated strong performance in producing semantically rich representations for long-document biomedical tasks, including clinical classification, retrieval, and entity recognition.
BioClinical ModernBERT further extended this line of work through continual pretraining on the largest biomedical and clinical corpus to date, comprising more than 53.5 billion tokens (Sounack2025BioClinicalModernBERT). Unlike earlier clinical encoders trained on narrow or single-institution datasets, this model uses data from multiple institutions, domains, and regions. This broader training data improves generalization and enables the model to achieve state-of-the-art performance across many biomedical NLP benchmarks.
Beyond English biomedical text, ModernBERT has also been shown to be effective in non-English and highly technical clinical settings. In Japanese radiology report classification, ModernBERT achieved a 39 per cent reduction in training time and up to a 1.651.65× speed-up during training and inference, while maintaining comparable classification performance (Yamagishi2025ModernBERTRadiology).
@@ 2.4.3 Architectural Analysis and Comparisons
In a controlled study designed to isolate architectural effects from differences in pretraining data, ModernBERT was compared with DeBERTaV3 when both models were pretrained on the same dataset. The results indicate that while ModernBERT offers faster training and inference speeds attributable to its architectural optimizations, the earlier DeBERTaV3 architecture maintained superior sample efficiency and overall benchmark performance under identical pretraining conditions (Antoun2025ModernBERT). This finding underscores the importance of disentangling architectural improvements from data effects when evaluating advances in encoder design and suggests that modern architectural enhancements may need to be paired with training methodology innovations to fully realize performance gains.
Another line of architectural comparison came from NeoBERT (Breton2025NeoBERT), a next-generation bidirectional encoder that incorporates state-of-the-art architectural refinements, modern pretraining regimes, and optimized depth-to-width design. NeoBERT achieves state-of-the-art results on large evaluation suites such as GLUE and the Massive Text Embedding Benchmark (MTEB), regardless of its relatively small model size of 250250M parameters and extended context length. It outperforms standard baselines, including BERT Large, RoBERTa Large, NomicBERT, and ModernBERT under identical fine-tuning conditions (Breton2025NeoBERT). This comparison highlights how architectural choices such as optimal layer configurations, positional encoding mechanisms, and more balanced model scaling can yield efficiency and performance advantages even compared to models with greater nominal capacity.
@@ 2.5 Turkish NLP Benchmarks
In response to these limitations, several Turkish NLP benchmark efforts have emerged that aim to consolidate multiple tasks under unified evaluation frameworks. Mukayese aggregated diverse Turkish NLP tasks, including language modeling, machine translation, NER, summarization, and text classification into a single benchmark suite, providing common baselines (Safaya2022Mukayese). TrGLUE adapted the GLUE paradigm to Turkish natural language understanding tasks, though its scope remained largely limited to sentence-level NLU (Altinok2025TrGLUE). Cetvel, a large-scale LLM-oriented evaluation benchmark, covered a broader set of discriminative and generative tasks while emphasizing cultural and linguistic relevance for Turkish (kuisai2024cetvel). More recently, TR-MTEB introduced a Turkish adaptation of the Massive Text Embedding Benchmark (MTEB), focusing on embedding and retrieval tasks (BaysanGungor2025TR-MTEB) by utilizing machine translation.
@@
3 TabiBERT: Model Architecture and Pretraining
In this section, we present a comprehensive overview of the pretraining procedure of the TabiBERT model, in terms of the pretraining dataset, tokenizer details, the pretraining objectives, and the model architecture.
3.1 Pretraining Data
We compiled a comprehensive dataset primarily consisting of Turkish monolingual text, supplemented with a small portion of English, coding and math data, to pretrain our model. To address significant size disparities and differences in quality across corpora and ensure balanced representation of different text types, we employed strategic oversampling of smaller but high-quality datasets. After applying the oversampling strategy, our final pretraining corpus consisted of 86.5886.58B tokens, of which $13$ per cent were non-Turkish sources. In the rest of the paper, all references to the size of the pretraining corpus will include the effects of oversampling. The details of each corpus are explained in the following subsections, and the training corpora statistics, as well as the sampling rates, are summarized in Table 2.
@@ 3.1 Pretraining Data
We extracted the tur_Latn subset of the FineWeb-2 (Penedo2025FineWeb2) as the main Turkish web corpus, with a knowledge cutoff date of April 2024. We also included a sample of 5 million English web pages from the FineWeb-2 English dataset to provide English language context during pretraining. Moreover, we included a Turkish Wikipedia dataset comprising articles scraped from Wikipedia’s Turkish-language pages in August 2024. This dataset provides encyclopedic knowledge covering a wide range of topics written in a formal and informative style.
@@
Table 2: Training Corpora Statistics: The complete pretraining corpus consists of 86.5886.58B tokens.
@@ 3.2 Tokenizer
Its design decisions were based on two recent works. Vocabulary size is selected to be computationally optimal for monolingual model training (Ali2024Tokenizer). Training algorithm and pre-tokenization RegEx are chosen based on their demonstrated effectiveness for LLM tokenization (Dagan2024Getting).
@@
The tokenizer was trained on a dataset mixture that consists of 95 per cent Turkish and 5 per cent English, code, and math, a similar mixture to our model’s pretraining data. The dataset mixture was curated to support a wide range of language capabilities, including proficiency in Turkish and English, as well as the ability to handle code snippets and mathematical expressions.
@@
Fertility (Rust2021HowGood) measures the average number of tokens required to represent text and serves as an indicator of tokenizer representation efficiency. While lower fertility is generally desirable in language modeling, vocabulary size and fertility involve clear trade-offs with model size and computational requirements (Ali2024Tokenizer).
@@
We measure fertility values by tokenizing 11M randomly selected web pages from our Turkish web corpus and report the relative fertility values compared to TabiBERT in Figure 1. We observe that:
@@
  • TabiBERT and BERTurk tokenizers, both being cased tokenizers, achieve comparable fertility values, despite TabiBERT having a significantly larger vocabulary size. This reflects TabiBERT tokenizer’s pre-tokenization RegEx design, which allocates dedicated tokens to digits, punctuation, and structural characters. This represents a conscious design choice that trades modest fertility increases for improved representation of structured content.
  • ytu-cosmos-BERT achieves the lowest fertility as a result of using an uncased tokenization strategy.
@@
  • As a result of being optimized for social media text, TurkishBERTweet has a similar fertility value to TabiBERT despite being uncased and having twice the vocabulary size.
  • Despite its large vocabulary of $256,000$ tokens, the mmBERT tokenizer exhibits significantly higher fertility than Turkish-specific tokenizers. This reflects the limited representation of Turkish in large-scale multilingual training corpora such as FineWeb-2 (Penedo2025FineWeb2), where the Turkish portion is only 0.67 per cent.333github.com/huggingface/fineweb-2/blob/main/fineweb2-language-distribution.csv
YTU-CosmosBERT(32K)BERTurk(32K)TurkishBERTweet(100K)mmBERT(256K)20-200202040409.4-9.42.6-2.6+1.2+1.2+41.0+41.0Relative Fertility (per cent)
Figure 1: Relative tokenizer efficiency, in per cents, on Turkish text across established Turkish BERT models and mmBERT, normalized to TabiBERT (dashed baseline). X-axis labels indicate models with their vocabulary sizes, while the Y-axis reports relative token counts required for text representation. Although TabiBERT’s conscious allocation of tokens to structural elements yields slightly higher fertility than BERTurk (comparable despite a larger vocabulary), this design choice enables superior handling of code, mathematical content, and document structure. The multilingual mmBERT, despite its 256K-token vocabulary, exhibits 41.0 per cent higher fertility, reflecting greater subword fragmentation on Turkish text.
3.3 Pretraining Objectives
For pretraining, we adopt the Masked Language Modeling (MLM) objective as in ModernBERT, with a 30 per cent token masking rate. We utilize the official ModernBERT pretraining codebase and recipe to ensure methodological consistency and reproducibility. The pretraining process is organized into 3 progressive phases, each exposing the model to 850B, 125B, and 25B tokens, respectively. Hence, we scale the ModernBERT pretraining recipe by 50 per cent with the aim of pretraining for 11T tokens in total.
To best utilize the dataset and maximize training throughput, we pre-tokenize the entire pretraining corpus and chunk it into pieces that fit within the context length. We perform optimization using the StableAdamW algorithm(Wortsman2023StableLowPrecision), paired with a trapezoidal learning rate schedule (Xing2018Walk) and 1-sqrt decay (Hagele2024Scaling) for improved stability during large-scale training. We pretrain our model for $117$ hours on a cluster of 8 x NVIDIA H100 GPUs, resulting in 11T tokens being exposed to the model, that is, $11.83$ epochs over the entire dataset.
3.4 Model Architecture
@@
Its inference speed is comparable to other models for a standard context length of 512; however, it is $2.65$ times faster than others when processing long contexts of $8,192$ tokens (Warner2024ModernBERT). Moreover, for both short and long context lengths, it can process inputs with twice the batch size thanks to its lower GPU memory footprint.
To facilitate replication and reproducibility, we release TabiBERT model weights on HuggingFace (TabiBERT-HF), with the hopes that it serves as a well-documented baseline incorporating ModernBERT’s architectural advances.
4 TabiBench: A Standardized Turkish NLP Benchmark
During the evaluation of TabiBERT, we encountered significant challenges in identifying appropriate fine-tuning datasets for establishing reliable baseline comparisons. A systematic review of commonly used Turkish NLP datasets revealed substantial fragmentation, which hampers the consistency and comparability of reported results. This fragmentation is also reflected in recent Turkish BERT studies, which frequently rely on ad hoc evaluation datasets or custom data splits tailored to individual works (Kesgin2023YTU; Najafi2023TurkishBERTweet).
Our analysis identified several recurring issues:
  • Quality inconsistencies: Several widely cited datasets suffer from inconsistent annotations, ambiguous labeling schemes, unclear annotation guidelines, or insufficient scale for robust evaluation.
  • Fragmented evaluation practices: Turkish NLP research lacks standardized train/validation/test splits. As a result, different studies adopt varying data subsets, preprocessing pipelines, and evaluation metrics for the same tasks, undermining the reliability of cross-study comparisons.
  • Limited task coverage: Existing benchmarks do not provide comprehensive coverage of the diverse task spectrum required to evaluate modern encoder-based models.
Such challenges are common in lower-resource language NLP but are rarely addressed systematically. Rather than compromise evaluation quality by using problematic datasets, we developed TabiBench as a standardized evaluation framework for Turkish NLP. TabiBench is designed around three core principles: